3. For each of the following productions determine whether there are IRS, CRS or DRS. a) y- ziz212
Answer #1

Answer 3

A function exhibit increasing returns to scale if we increase all inputs by t times output gets increased by more than t times, A function exhibit decreasing returns to scale if we increase all inputs by t times output gets increased by less than t times and A function exhibit constant returns to scale if we increase all inputs by t times output gets increased by exact t times. Note that t must be greater than 1.

Mathematically a function exhibit increasing returns to scale(IRS) if F(tx , ty) > tF(x,y).

A function exhibit Decreasing returns to scale(DRS) if F(tx , ty) < tF(x,y)

a function exhibit Constant returns to scale(CRS) if F(tx , ty) = tF(x,y)

a)

y = f(z1 , z2) = z1z21/2

f(tz1 , tz2) = t3/2z1z21/2 > tz1z21/2 = tf(z1 , z2) for all t > 1

Hence this function exhibit increasing returns to scale(IRS)

b)

y = f(z1 , z2) = z12 + z2

f(tz1 , tz2) = t2z12 + tz2 = t(tz12 + z2) > t(z12 + z2) = tf(z1 , z2) for all t > 1

Hence this function exhibit increasing returns to scale(IRS)

(c)

y = f(z1 , z2) = 3z1 + 2z21/2  

f(tz1 , tz2) = 3tz1 + 2t1/2z21/2 = t(3z1 + 2(z2/t)1/2) < t(3z1 + 2z21/2) = tf(z1 , z2) for all t > 1

Hence this function exhibit decreasing returns to scale(DRS)